OpenStudy (anonymous):

Find side c if side a = 10 inches and side b = 4 inches.

OpenStudy (anonymous):

Well... are you given that ABC is a right triangle, and that the c, or the missing side, is the hypotenuse?

OpenStudy (anonymous):

^ I'm assuming yes, that's the case

OpenStudy (texaschic101):

use pythagorean theorem

OpenStudy (anonymous):

\[10^{2}+4^2 = 116\]

OpenStudy (anonymous):

\[\sqrt{116}=10.77\]

OpenStudy (anonymous):

c= 10.77

OpenStudy (anonymous):

\[\sqrt{10^2+4^2}=\sqrt{116}=2 \sqrt{29}≈10.77\] that is if we assume that c is the hyp., .... but that is not necessarily the case. \[\sqrt{10^2-4^2}=\sqrt{84}=2 \sqrt{21}≈9.17\] can as well be though that 10 is the hyp., but i think that in your case, the hyp, is the MISSING side C.

OpenStudy (texaschic101):

a^2 + b^2 = c^2 10^2 + 4^2 = c^2 100 + 16 = c^2 116 = c^2 sq rt 116 = c 10.77 = c

OpenStudy (texaschic101):

thats only if c is the hypotenuse